OpenStudy (anonymous):

Ronald rode away from home in a car at the rate of 32mph . he walked back at the rate of 4 mph. the round trip required 2 hours and 15 minutes. how far did Ronald ride?

OpenStudy (inkyvoyd):

Comon, we just did one of these for you!

OpenStudy (anonymous):

same idea as before

OpenStudy (saifoo.khan):

it's very similar.

OpenStudy (saifoo.khan):

Did you tried the question @Winnie846039137

OpenStudy (anonymous):

expression for time going is \(\frac{d}{32}\) and expression for time returning is \(\frac{d}{4}\) total time is 2 hours and 15 minutes aka \(2\tfrac{1}{4}\) hours set \[\frac{d}{32}+\frac{d}{4}=\frac{9}{4}\] and solve for d

OpenStudy (mathteacher1729):

This is the FIRST idea which will take some "meditation time": \[ \Huge \text{Distance } = \text{ rate} \times \text{ time} \] If you are waking at 4 miles per hour that means 4 miles = 1 hr 8 miles = 2 hrs 12 miles = 3 hrs etc. If you are driving 32 miles per hour that means 32 miles = 1 hr 64 miles = 2 hrs 96 miles = 3 hrs etc. This is the SECOND idea which will take "meditation time": You can find out how long it takes you to drive any amount of miles you like, by dividing both sides by the number of miles. 4 miles = 1 hour means 1 mile = 1/4 hours 32 miles = 1 hour means 1 mile = 1/32 hours You need to understand BOTH Of these ideas before the solutions given will make sense.

OpenStudy (anonymous):

one thing to be careful of is that the units you are given is miles and hours, so don't mess with the minutes you want to write everything in terms of hours, which is why i wrote \[\frac{9}{4}\] hours